(Appendix 4B) Sequential Method of Support Department Cost Allocation

Stevenson Company is divided into two operating divisions: Battery and Small Motors. The company allocates power and general factory costs to each operating division using the sequential method. General Factory is allocated first in the sequential method for the company. Support department cost allocations using the sequential method are based on the following data:

Support Departments Operating Divisions

Power
General Factory
Battery
Small Motors
Overhead costs $160,000 $430,000 $163,000 $84,600
Machine hours 2,000 2,000 8,000 1,000
Square footage 1,000 1,500 5,000 20,000
Direct labor hours 18,000 60,000

Required:

1. Calculate the allocation ratios for Power and General Factory. (Note: Carry these calculations out to four decimal places.)

Allocation ratios for General Factory

Power
Battery
Small Motors

Allocation ratios for Power

Battery
Small Motors

2. Allocate the support service costs to the operating divisions. Round all amounts to the nearest dollar. Use a minus sign to indicate a subtraction. For those boxes in which no entry is required, leave the box blank or enter zero ("0").

Support Departments Operating Divisions

Power

General Factory

Battery
Small
Motors
Direct costs $ $ $ $
Allocate:
    General Factory
    Power
Total $ $ $ $

3. Assume divisional overhead rates are based on direct labor hours. Calculate the overhead rate for the Battery Division and for the Small Motors Division. Round overhead rates to the nearest cent.

Battery overhead rate $
Small Motors overhead rate $
Answer #1

1. Calculate the allocation ratios for Power and General Factory. (Note: Carry these calculations out to four decimal places.)

Allocation ratios for General Factory

Power (1000/26000) 0.0385
Battery (5000/26000) 0.1923
Small Motors (20000/26000) 0.7692

Total square footage= 1000+5000+20000= 26000

Allocation ratios for Power

Battery (8000/9000) 0.8889
Small Motors (1000/9000) 0.1111

Total machine hours of the operating departments= 8000+1000= 9000

2. Allocate the support service costs to the operating divisions. Round all amounts to the nearest dollar. Use a minus sign to indicate a subtraction. For those boxes in which no entry is required, leave the box blank or enter zero ("0").

Support Departments Operating Divisions

Power

General Factory

Battery
Small
Motors
Direct costs $ 160000 $ 430000 $ 163000 $ 84600
Allocate:
    General Factory 16555 -430000 82689 330756
    Power -176555 156940 19615
Total $ 0 $ 0 $ 402629 $ 434971

Allocation of General Factory department

Power= 0.0385*$430000= $16555

Battery= 0.1923*$430000= $82689

Small motors= 0.7692*$430000= $330756

Allocation of Power department

Total cost of Power department= $160000+16555= $176555

Battery= 0.8889*$176555= $156940

Small motors= 0.1111*$176555= $19615

3. Assume divisional overhead rates are based on direct labor hours. Calculate the overhead rate for the Battery Division and for the Small Motors Division. Round overhead rates to the nearest cent.

Battery overhead rate ($402629/18000) $22.37
Small Motors overhead rate ($434971/60000) $7.25